# Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese

### v1

@@ 5.1.3 Options and Answer Analysis
Options Quality
We also evaluate the options in a similar manner to the evaluation of the question. From Table 6, we observe that LLM_Gen generates higher quality options compared to LLM_Adapt. However, there is still a significant gap in performance between Indonesian and Sundanese. This once again highlights the performance discrepancy between medium- and lower-resource languages.
Common Mistakes
We conduct a detailed analysis to identify the common errors found in each dataset, shown in Table 7. For LLM_Adapt, the most common issues are typos and mechanical errors, particularly with capitalization (e.g., "indonesia" instead of "Indonesia"), along with translation errors. In the Sundanese dataset, the errors are mainly due to major translation issues. As for LLM_Gen, although errors constitute less than 10% of the total evaluated samples, the primary issue is the presence of invalid options, particularly when no correct answers are among the options. For Sundanese, alongside translation errors, the model also produces some invalid options.
@@
Error Type % of error type
@@
Translation 33.33% 48.51% 0.00% 13.61%
Sent. structure 12.82% 11.94% 0.00% 0.00%
Sent. fluency 28.21% 27.61% 36.17% 36.73%
Sent. context 7.69% 2.24% 9.57% 16.33%
Subjectivity 2.56% 0.75% 0.00% 0.00%
Typo/mechanics 15.38% 8.96% 54.26% 33.33%
@@
Dataset % of correct choices
@@
Table 6: Choices generation accuracy of LLM-generated datasets. The accuracy reflects the percentage of ‘clean’ choices.
Error Type % of error type
@@
Translation 10.17% 77.32% 0.00% 76.61%
Sent. fluency 1.69% 1.03% 9.52% 3.23%
Sent. context 1.69% 1.03% 0.00% 0.00%
Invalid options 3.39% 2.06% 80.95% 20.16%
Typo/mechanics 83.05% 18.56% 9.52% 0.00%
Table 7: Distribution of choices generation error types of LLM-generated datasets.
5.2 LLM vs. Human-Generated Data
To compare the data generated by LLM with human-generated data, we analyze the proportion of shared unigram and bigram tokens between the LLM_Gen and Human_Gen datasets. The overlap percentage is calculated by dividing the number of shared tokens by the total number of tokens. We find that the unigram overlap percentage is 39.75% for Human_Gen and 65.48% for LLM_Gen. A similar trend is observed for the bigram overlap percentage, with 12.41% for Human_Gen and 15.98% for LLM_Gen. This shows that many tokens present in Human_Gen also exist in LLM_Gen, but the reverse is not equally true. Given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol. [^4] In the context of culinary, ‘kalis’ means a state of dough that is well-kneaded and ready to be processed further. ‘Cimol’ is an Indonesian street food made from tapioca flour.
Upon sample-level analysis, we also find token variance disparities, particularly related to cultural context. For instance, given ‘kerupuk’ (crackers) concept, human annotators can formulate questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers). In contrast, LLM tends to ask more general questions, like "What is the common primary ingredient of crackers?" Also, although LLM can generate some culturally related entities such as ‘tari kecak’, a traditional dance from Bali, it is more popular compared to ‘tari pendet’, which is generated by the human annotator. More examples are outlined in Table 11.
@@ 6.1 Experiment Setup
We include PolyLM-13B Wei et al. (2023), an LLM trained on a multilingual dataset, predominantly in English and Chinese; BLOOMZ-7B Muennighoff et al. (2022), an LLM fine-tuned on the xP3 dataset; and SeaLLM-7B Nguyen et al. (2023), an LLM with a focus on Southeast Asian languages.
@@
We include Merak-v4 Ichsan (2023), an Indonesian LLM fine-tuned based on MistralOrca-7B. We also include MalaysianMistral-7B Zolkepli et al. (2024), a Malaysian LLM extended from Mistral-7B.
@@
We include GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo, OpenAI (2023), which are accessed via OpenAI API.
@@ 6.2 Overall Performance
Figure 3: Overall performance of LLMs on our combined test set. We run three different prompts for each question in a zero-shot setting.
@@
Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen.
To assess LLMs’ ability to answer both LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets. Our result, shown in Figure 4, reveals that LLMs perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data. This indicates that the model may possess some knowledge to answer LLM-generated questions, even though it was not specifically trained on Indonesian or Sundanese data. However, the model still struggles to answer new knowledge present in Human_Gen data. The performance gap is also observed in Indonesian LLM, Merak-v4. The accuracy of this model in Sundanese is also low, indicating that training LLM solely on Indonesian texts is not enough to make it perform well in local languages such as Sundanese.
@@ 6.4 Performance by Question Category
Figure 5 shows the LLMs’ performance across different question categories. It reveals that LLMs perform better in the activity and place questions, but struggle in culinary. A notable performance gap exists between LLM_Gen and Human_Gen dataset, with weaker results on the latter. Moreover, open LLMs exhibit a significant performance gap compared to proprietary models, especially in Human_Gen, indicating their shortcomings in capturing commonsense knowledge. Interestingly, a significant improvement is seen from GPT-4 (gpt-4-0613) to GPT-4 Turbo (gpt-4-1106), especially in Sundanese Human_Gen. This suggests that the model is acquiring more knowledge, one reason for which possibly comes from the interactions on the ChatGPT web. Despite these advancements, some categories still score below 80%, showing there is still room for improvement.
@@ 7 Discussion
Since our dataset is in a multiple-choice format, LLMs might look better than they actually are because they can just pick one answer from the given options. To see LLMs’ genuine capability, we test GPT-4 Turbo on 100 randomly sampled questions from Indonesian Human_Gen in open-ended settings, i.e., asking the question directly without showing answer options or providing any extra instructions. Our manual evaluation shows a 77% accuracy rate, highlighting the model’s struggle with open-ended questions and confirming our dataset’s importance. Among the wrong answers, 18 cases stemmed from overly general answers not specific to Indonesia. For example, when asked "Where can we find many villas?" the model answers "Bali" but also listed other cities in mostly Western countries. In contrast, the annotator writes "Bandung" as the correct answer. As for the other wrong cases, the model struggles the most in answering questions in the culinary category. More failure examples can be seen in Table 12.
7.2 Extension to Other Local Languages
Besides Indonesian, our study focuses on one Indonesian local language, Sundanese. As previously discussed, Indonesia has many local languages; however, we cannot cover all of them due to resource constraints. We aim for our findings in Sundanese to act as a starting point for other languages. We anticipate that LLMs might perform worse than Sundanese for extremely low-resource languages, like Buginese or Toba Batak. Javanese, on the other hand, is expected to have comparable performance to Sundanese Winata et al. (2023); Bang et al. (2023). Generating good-quality synthetic data could be particularly beneficial for extremely low-resource languages, such as Buginese or Toba Batak, given the difficulty of finding native speakers. Yet, our case study in Sundanese indicates that LLMs might struggle even more with understanding cultural nuances in such languages.
@@ 8 Conclusion
In this study, we created Indonesian and Sundanese CommonsenseQA datasets using various dataset collection methods, including automatic dataset generation utilizing LLMs and manual dataset generation employing human annotators from diverse regions. Our thorough analysis reveals a substantial gap in the performance of LLMs between Indonesian and Sundanese, both in the dataset generation and in answering the questions in the dataset itself. Thus, one must be cautious when using existing LLMs in generating datasets, particularly for lower-resource languages, since their culturally relevant commonsense knowledge may be limited. Our findings highlight the urgent need to create high-quality resources for these underrepresented languages to prevent them from being left behind.
@@
Our work has some limitations that should be acknowledged. Firstly, in terms of language coverage, we were only able to cover Indonesian and Sundanese due to the available resources and the authors’ familiarity with these languages. Additionally, the annotators we recruited were mostly from Java island, with one annotator from Bali island. Despite our effort to include a range of question concepts from different regions, including those beyond Java and Bali islands, it is possible that some bias may exist, especially in the Indonesian dataset. This is because the questions were generated primarily by annotators from Java and Bali, and their perspectives and cultural backgrounds may have influenced the content. Nonetheless, we have taken measures to eliminate potentially harmful questions.
Additionally, we use a multiple choice question format, following English CommonsenseQA data format Talmor et al. (2019) to facilitate a more straightforward and robust evaluation process. Although open-ended generation may offer a more challenging benchmark for LLMs, as discussed in Section 7.1, evaluating LLMs in such settings poses its own set of challenges, especially in low-resource languages where ‘LLM-as-a-judge’ approach may not be as effective as in English. Nevertheless, our dataset can serve as a starting point for this line of research direction.
@@ A.1.2 Automatic Data Generation
Table 8 shows the win rate accuracy of concept, question, and choices generation of Indonesian LLM, Merak-v4, and best-performing proprietary LLM, GPT-4 Turbo. The results demonstrate that GPT-4 Turbo significantly outperforms Merak-v4, with win rates ranging from 72% to 96%. Our sample-level analysis indicates that a significant number of questions generated by Merak-v4 tend to have obvious answers or are formulated as yes/no questions, such as "Apakah jagung dapat dimakan?" (eng: Is corn edible?). Additionally, despite being trained on Indonesian texts, the model occasionally produces questions with US-centric knowledge. For instance, it generates "Siapa yang memimpin kampanye pemilihan presiden pertama di Amerika Serikat?" (eng: Who led the first presidential election campaign in the United States?).
A.3 Additional Analysis of the Generated Data: Common Mistakes in Concept Adaptation
@@
eng normal-$\rightarrow$ ind ind normal-$\rightarrow$ sun
@@
Appendix B Human Annotation Guideline
@@
Appendix C Annotators Demographics
Data Team Ethnicity Domicile Num
Indonesian Sundanese Sukabumi, West Java 1
Depok, West Java 1
Javanese Magelang, Central Java 1
Bojonegoro, East Java 1
Betawi Tangerang, Banten 1
Balinese Denpasar, Bali 1
Sundanese Sundanese Bandung, West Java 2
Bogor, West Java 1
Majalengka, West Java 1
Sukabumi, West Java 2
Total 12
In accordance with the aim of constructing the dataset, we involved all Indonesian native-speaker annotators in the manual data generation process. All the annotators who worked on the Indonesian dataset were people from several regions on the islands of Java and Bali. Meanwhile, the annotators involved in building the Sundanese dataset were Sundanese people who come from several different regions in West Java. Each dataset development team consists of 6 annotators. The detailed annotators’ demographics are shown in Table 10.
@@ Appendix D Manual Data Generation Findings
During the manual data generation process, we discovered several interesting findings from the data created by the annotators.
@@
We found quite a lot of interesting findings in the process of creating commonsense questions. We found that there were questions created using very general contexts, while the limitations of this data generation were the Indonesian and Sundanese contexts (specifically for the Sundanese dataset). For instance, "Apa jenis restoran yang paling terkenal di seluruh dunia?" (eng: Which type of restaurant is most famous worldwide?) In that question, the annotator developed the concept question “restoran” (eng: restaurant) into a question with a very general context: “di seluruh dunia" (eng: all over the world). We also found subjective questions, so the answers that emerged also had subjective value (not commonsense). These subjective questions are usually characterized by the use of superlative adjectives. Another type of finding related to the question category is the creation of logical/causal questions. This seems to be based on confusion from annotators regarding the boundaries of commonsense and logic/causation.
Choices
The findings related to choices that are most often encountered are making choices that have the same value as each other so that it is difficult to determine the answer and distractor. For instance, the question "Kue apa yang biasanya disajikan pada momen lebaran?" (eng: What cakes are usually served during Eid?). All the options given are types of cakes that are generally served during Eid. This finding is also related to the issue of subjectivity.
@@
Answers and distractors are another category that also has many findings in the manual data generation process. This issue is not much different from the findings in Choices: answers and distractors are equivalent. Apart from that, in this category, it was also found that personal experience was used in determining answers so that more common answers were determined as distractors. For instance, the question "Apa yang biasanya orang lakukan di stasiun?" (eng: What do people usually do at the train station?). The annotator specified "mengantar teman/kerabat" (eng: accompany friends/relatives) as the answer. While the "naik-turun kereta" (eng: get on and off the train) option was chosen as a distractor.
@@
As we employ annotators from different regions, it enables us to explore variations in the data they generate. We examine this by calculating the number of answer conflicts that arise during the "answering question" phase of our data generation pipeline (§4.2). The results from Indonesian and Sundanese annotators are detailed in Figure 6 and 7, respectively. Our analysis revealed that, out of the Indonesian annotators, the one from Bali has the highest number of conflicting answers. However, the number of questions generated by the Bali annotator does not seem to have a lot of conflicts, suggesting that the annotator tends to generate easier questions, some of which have an obvious answer. This finding contrasts with the results from West Java annotators, who, despite generating questions that lead to a higher number of conflicts, do so mainly due to the creation of more challenging option distractors.

### v2

@@ 5.1.3 Options and Answer Analysis
Error Type % of questions
@@
No error 75.32% 15.19% 68.67% 51.00%
[…]
Sent. structure 3.16% 10.13% 0.00% 0.00%
Sent. fluency 6.96% 23.42% 11.33% 18.00%
Sent. context 1.90% 1.90% 3.00% 8.00%
Subjectivity 0.63% 0.63% 0.00% 0.00%
Typo/mechanics 3.80% 7.59% 17.00% 16.33%
@@
Dataset % of correct options
@@
Table 5: Options generation accuracy of LLM-generated datasets, measured by the % of clean/error-free options.
Error Type % of options
@@
No error 62.66% 38.61% 93.00% 58.67%
[…]
Sent. fluency 0.63% 0.63% 0.67% 1.33%
Sent. context 0.63% 0.63% 0.00% 0.00%
Invalid options 1.27% 1.27% 5.67% 8.33%
Typo/mechanics 31.01% 11.39% 0.67% 0.00%
Table 6: Distribution of options generation error types of LLM-generated datasets.
Options Quality
We also evaluate the options in a similar method to the evaluation of the question. From Table 5, we observe that LLM_Gen generates higher quality options compared to LLM_Adapt. However, there is still a significant gap in performance between Indonesian and Sundanese, once again highlighting the performance discrepancy between medium- and lower-resource languages.
Common Mistakes
As detailed in Table 5.1.3, for LLM_Adapt, the most common issues are minor typos and mechanical errors, particularly with capitalization (e.g., "indonesia" instead of "Indonesia"). For Sundanese, the errors are mainly due to major translation errors. As for LLM_Gen, the most common errors involve the presence of invalid options, particularly when no correct answers are among the options. For Sundanese, alongside translation errors, the model also produces some invalid options. However, it is important to note that the proportion of invalid options is very low for both languages, indicating that the model is still capable of generating a large number of valid options.
5.2 LLM vs. Human: Lexical Diversity
To compare the lexical diversity between LLM and human-generated data, we analyze the proportion of shared tokens between the LLM_Gen and Human_Gen, calculated by dividing the number of unique shared tokens by the total number of unique tokens. We find that the unigram overlap percentage is 39.75% for Human_Gen and 65.48% for LLM_Gen. A similar trend is observed for the bigram overlap percentage, with 12.41% for Human_Gen and 15.98% for LLM_Gen. This shows that many tokens present in Human_Gen also exist in LLM_Gen, but the reverse is not equally true. We also find that Human_Gen dataset has 8,596 unique tokens, higher than LLM_Gen with 6,677 unique tokens.
Upon some sample-level analysis, we find that given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol. [^4] In the context of culinary, ‘kalis’ means a state of dough that is well-kneaded and ready to be processed further. ‘Cimol’ is an Indonesian street food made from tapioca flour. Although LLM can still generate some valid concepts, it tends to ask more general questions or popular concepts. For instance, given ‘kerupuk’ (crackers) concept, human annotators can formulate questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers). In contrast, LLM tends to ask more general (but still relevant) questions, like "What is the common primary ingredient of crackers?" More examples are shown in Table 12.
@@ 6.1 Experiment Setup
We include PolyLM-13B Wei et al. (2023), which trained on a multilingual dataset (mostly English and Chinese); BLOOMZ-7B Muennighoff et al. (2022), which fine-tuned on xP3; SeaLLM-7B Nguyen et al. (2023), which covers Southeast Asian languages; and Aya Üstün et al. (2024), recent open LLM trained on 101 languages, including Indonesian and Sundanese.
@@
We include Merak-v4 Ichsan (2023), an Indonesian LLM based on MistralOrca and MalayMistral-7B Zolkepli et al. (2024), a Malaysian LLM extended from Mistral.
@@
We include GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo OpenAI (2023).
@@
Figure 3: Overall performance of LLMs on our combined test set.
Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen dataset in Indonesian and Sundanese.
@@ 6.2 Overall Performance
We also observe a substantial gap between Indonesian and Sundanese (10–20% accuracy drop), suggesting that current LLMs struggle to answer our Sundanese commonsense questions, even in multiple-choice settings. The huge gap, particularly in Merak-v4, highlights the limitations of training LLMs solely on Indonesian texts, as it does not ensure transferable performance across other local languages, especially due to the morphological differences between the languages. Including Sundanese texts in the training data, as demonstrated by improved performance in the Aya model, has proven effective in narrowing the performance gap between Indonesian and Sundanese.
@@
To assess LLMs’ ability to answer LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets. Figure 4 shows that LLMs generally perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data. This indicates that the model may possess some knowledge to answer LLM_Gen questions, even though it was not specifically trained on Indonesian or Sundanese data. Still, it is important to note that even in LLM_Gen, MistalOrca’s average score is still lower compared to LLMs that include Indonesian texts in their training.
Interestingly, as the model gets stronger, we observe that the gap tends to narrow, especially in GPT models. For Aya, despite its relatively stronger performance compared to other open-source LLMs, its performance is significantly dropped in Human_Gen data. This decline is likely due to Aya being a ‘translation-heavy’ model, with a large proportion of its training data derived from translations rather than human annotations.
@@ 6.4 Performance by Question Category
Figure 8 shows the LLMs’ performance across different question categories. It shows that LLMs perform better in the activity and place questions, which benefit from a lot of information readily available on the internet. However, they struggle with culinary questions, where specialized or cultural knowledge is often required. Interestingly, a significant improvement is seen from GPT-4 (0613 ver.) to GPT-4 Turbo (1106 ver.), particularly in Sundanese Human_Gen. This suggests that the model is acquiring more knowledge, which possibly comes from the interactions on the ChatGPT web. However, some categories still score below 80%, showing there is still room for improvement.
@@ 7 Discussion
[…]
Among the wrong answers, most cases come from overly general answers not specific to Indonesia, especially for questions in culinary, culture, and activity categories. For instance, given unique activity questions such as: "What is the mandatory song to be sung during the ‘mengheningkan cipta’ (moment of silence) in the flag ceremony?" the model gives a wrong answer in open-ended generation settings even if it could identify the correct answer in the multiple-choice setup. More failure examples can be seen in Table 13.
7.2 LLM-generated Data Quality
Due to the high performance of LLMs in answering LLM-generated data, one might argue that the data is less valuable. Ideally, creating a larger human-generated dataset is the most straightforward method to create a high-quality dataset. However, this approach may not always be feasible in the context of underrepresented languages, particularly given the limited resources. Building a dataset manually can be costly and challenging (we spent more than \$3,000 in total to build Human_Gen), and leveraging LLMs could be a practical and cheaper alternative, especially if we want to scale up the data.
To address the limitation of LLM-generated data, we can apply additional processes to reduce the noise. For instance, a collaboration between humans and models can be conducted to fix and revise the synthetic questions Liu et al. (2022); Putri and Oh (2022). This noise can also be reduced automatically, especially in Indonesian LLM_Gen questions, where errors typically involve minor typos. It is also important to note that not all LLM-generated data have low quality. Our experiments demonstrate that LLM_Gen has significantly fewer errors compared to LLM_Adapt, particularly in Sundanese. We also compared the performance between the cleaned vs. raw version of this data (Appendix F) and found that the performance variance between the original and cleaned data of LLM_Gen is relatively minor.
@@ 8 Conclusion
In this study, we introduced new Indonesian and Sundanese CommonsenseQA datasets built using various methods, including automatic dataset adaptation and generation with LLMs, as well as manual dataset creation by human annotators from diverse regions. Our thorough analysis reveals that adaptation from the existing English data is less effective for Sundanese. In contrast, when directly generated in the target languages, GPT-4 Turbo can produce relevant knowledge in both languages, although the cultural depth has not yet matched the human-generated ones. Nevertheless, given the limited resources, combining LLMs and humans for dataset creation can be a practical solution, allowing for a more efficient dataset creation by reducing the over-reliance on a single data source.
Our goal is for LLMs to serve as beneficial tools for diverse communities, not solely those from higher-resource languages, cultures, or regions. This is especially important since LLMs are continuously being used to generate data. We hope our work represents a crucial step toward achieving this broader objective.
@@
Language and Region Coverage
In terms of language coverage, we were only able to cover Indonesian and Sundanese due to the available resources and the authors’ familiarity with these languages. Additionally, the annotators we recruited were mostly from Java island, with one annotator from Bali island. Despite our effort to include a range of question concepts from different regions, including those beyond Java and Bali islands, it is possible that some bias may exist, especially in the Indonesian dataset. This is because the questions were generated primarily by annotators from Java and Bali, and their perspectives and cultural backgrounds may have influenced the content. Nonetheless, we have taken measures to eliminate potentially harmful or stereotypical questions.
Extension to Other Local Languages
Besides Indonesian, our study focuses on one Indonesian local language, Sundanese. As previously discussed, Indonesia has many local languages; however, we cannot cover all of them due to resource constraints. We aim for our findings in Sundanese to act as a starting point for other languages. We anticipate that LLMs might perform worse than Sundanese for extremely low-resource languages, like Buginese or Toba Batak. Javanese, on the other hand, is expected to have comparable performance to Sundanese Winata et al. (2023); Bang et al. (2023). Generating synthetic data could be particularly beneficial for extremely low-resource languages, given the difficulty of finding native speakers. Yet, our case study in Sundanese indicates that LLMs might struggle even more with generating fluent sentences and/or culturally unique concepts in such languages. Still, we believe that the combination or even collaboration between LLM and humans may serve as a starting point.
Multiple-Choice Format
We use a multiple choice question format, following English CommonsenseQA data format Talmor et al. (2019) to facilitate a more straightforward and robust evaluation process. Although open-ended generation may offer a more challenging benchmark for LLMs, as also discussed in Section 7.1, evaluating LLMs in such settings poses its own set of challenges, especially in low-resource languages where ‘LLM-as-a-judge’ approach may not be as effective as in English. Nevertheless, our dataset can serve as a starting point for this line of research direction.
@@ A.1.2 Automatic Data Generation
Table 7 shows the win rate accuracy of concept, question, and options generation of Indonesian LLM, Merak-v4, and best-performing proprietary LLM, GPT-4 Turbo. The results demonstrate that GPT-4 Turbo significantly outperforms Merak-v4, with win rates ranging from 72% to 96%. Our sample-level analysis indicates that a significant number of questions generated by Merak-v4 tend to have obvious answers or are formulated as yes/no questions, such as "Apakah jagung dapat dimakan?" (eng: Is corn edible?). Additionally, despite being trained on Indonesian texts, the model occasionally produces questions with US-centric knowledge. For instance, it generates "Siapa yang memimpin kampanye pemilihan presiden pertama di Amerika Serikat?" (eng: Who led the first presidential election campaign in the United States?).
A.3 Additional Analysis of the LLM-generated Data: Common Mistakes in Concept Adaptation
Status Num (%) Concept Example
Orig (ind) Modified (sun)
Correct 122 (77.22%)
hujan abu vulkanik
(volcanic ashfall)
hujan lebu vulkanik
(volcanic ashfall)
menara (tower) munara (tower)
Wrong 36 (22.78%)
cicak
(house gecko)
kadal imah
(house lizard)
klinik gigi
(dental clinic)
klinik dental
(dental clinic)
@@ A.3 Additional Analysis of the LLM-generated Data: Common Mistakes in Concept Adaptation
eng $\rightarrow$ ind ind $\rightarrow$ sun
@@
Appendix B Annotators Demographics
Data Team Ethnicity Domicile Num
Indonesian Sundanese Sukabumi, West Java 1
Depok, West Java 1
Javanese Magelang, Central Java 1
Bojonegoro, East Java 1
Betawi Tangerang, Banten 1
Balinese Denpasar, Bali 1
Sundanese Sundanese Bandung, West Java 2
Bogor, West Java 1
Majalengka, West Java 1
Sukabumi, West Java 2
Total 12
In accordance with the aim of constructing the dataset, we involved 6 Indonesian and 6 Sundanese native-speaker annotators. We chose annotators who have prior experience in building NLP datasets in Indonesian and/or Sundanese through close recruitment. All the annotators who worked on the Indonesian dataset were people from several regions on the islands of Java and Bali. Meanwhile, the annotators involved in building the Sundanese dataset were Sundanese people who come from several different regions in West Java. The detailed annotators’ demographics are shown in Table 10.
Appendix C Human Annotation Guideline
@@ Appendix D Manual Data Generation Findings
During the manual data generation process, we discovered several interesting findings from the data created by the annotators. Note that the common errors within these findings occur before the quality control meeting (§4.2). In our final dataset, we have ensured that such problems have been fixed.
@@
We found quite a lot of interesting findings in the process of creating commonsense questions. We found that there were questions created using very general contexts, while we required the data generation to be scoped within the Indonesian and Sundanese contexts (specifically for the Sundanese dataset). For instance, an annotator produced "Apa jenis restoran yang paling terkenal di seluruh dunia?" (eng: Which type of restaurant is most famous worldwide?) In that question, the annotator developed the concept question “restoran” (eng: restaurant) into a question with a very general context: “di seluruh dunia" (eng: all over the world). We also found subjective questions, so the answers that emerged also had subjective value (not commonsense). These subjective questions are usually characterized by the use of superlative adjectives. Another type of finding related to the question category is the creation of logical/causal questions. This seems to be based on confusion from annotators regarding the boundaries of commonsense and logic/causation.
Options
The findings related to options that are most often encountered are making choices that have the same value as each other so that it is difficult to determine the answer and distractor. For instance, the question "Kue apa yang biasanya disajikan pada momen lebaran?" (eng: What cakes are usually served during Eid?). All the options given are types of cakes that are generally served during Eid. This finding is also related to the issue of subjectivity.
@@
Answers and distractors are another category that also has many findings in the manual data generation process. This issue is not much different from the findings in Options: answers and distractors are equivalent. Apart from that, in this category, it was also found that personal experience was used in determining answers so that more common answers were determined as distractors. For instance, the question "Apa yang biasanya orang lakukan di stasiun?" (eng: What do people usually do at the train station?). The annotator specified "mengantar teman/kerabat" (eng: accompany friends/relatives) as the answer. While the "naik-turun kereta" (eng: get on and off the train) option was chosen as a distractor.
@@
As we employ annotators from different regions, it enables us to explore variations in the data they generate. We examine this by calculating the number of answer conflicts that arise during the "answering question" phase of our data generation pipeline (§4.2). The results from Indonesian and Sundanese annotators are detailed in Figure 5 and 6, respectively. Our analysis revealed that, out of the Indonesian annotators, the one from Bali has the highest number of conflicting answers. However, the number of questions generated by the Bali annotator does not seem to have a lot of conflicts, suggesting that the annotator tends to generate easier questions, some of which have an obvious answer. This finding contrasts with the results from West Java annotators, who, despite generating questions that lead to a higher number of conflicts, do so mainly due to the creation of more challenging option distractors.